What is the change in internal energy that if a system that absorbs 244 J of heat and does 215 J of work?


1
Expert's answer
2022-05-23T16:05:54-0400
"E=U+W"

"U=E-W"

"U=244J-215J=29J"

Internal energy increases by 29J.